OpenStudy (anonymous):

Express in Sigma Notation: 2+1+1/2+.....+1/128

OpenStudy (anonymous):

I know it is a Geometric Series, and i know that to find the nth power is use the formula a_n=a_1(r)^n-1, but i dont know how to solve it after i plug in the values.

OpenStudy (anonymous):

\[\sum_{i=1}^{9} 2 (1/2)^{n-1}\]

OpenStudy (anonymous):

what is the limit you put above the sigma?

OpenStudy (anonymous):

the reason i put the formula in my first comment is because i was trying to find a_n(the"place" of 1/128) is so that i know at what number place the series ends.

OpenStudy (anonymous):

AKA the number above the sigma symbol

OpenStudy (anonymous):

upper limit is 9

OpenStudy (zzr0ck3r):

I wouldn't call it a limit

OpenStudy (anonymous):

and did you use the formula\[a _{n}=a _{1}\times(r )^{n-1}\]

OpenStudy (anonymous):

well not the limit, but the last number of that series

OpenStudy (zzr0ck3r):

correct, the largest index

OpenStudy (anonymous):

yesh. well could someone run me through the process of solving for "n" or the largest index? ......i dont know how to solve because "n" is in the exponent.....

OpenStudy (zzr0ck3r):

um you are not solving for n, n is just a counter to let you know you will stop after n = 9

OpenStudy (anonymous):

what i mean is how do i find our that n=9?

OpenStudy (zzr0ck3r):

\(\sum_{i=0}^3 i = 0+1+2+3 = 6\)

OpenStudy (zzr0ck3r):

\(\sum_{i=1}^{9} 2 (1/2)^{n-1}=2 (1/2)^0+2 (1/2)^1+2 (1/2)^2+2 (1/2)^3+......+2 (1/2)^9\)

OpenStudy (anonymous):

so you just kept solving the series until you got to 1/128? see what if the largest index was 100? It would be impractical to keep solving until i reach the value at 100...so thats why i was using the Geometric Series sum formula to solve for the place or the highest index which in the formula is "n"

OpenStudy (anonymous):

the question you posted was only asking to express in sigma notation not find the nth term yes, if you had an index up to 100, it will be better to use the formula that solves for the nth term, but in this case you just asked to express it in sigma notation

OpenStudy (anonymous):

but the thing is the last number is 1/128, but it only tells me the number, not the actually place which is apparently 9. Thats why i was using the formula to find "n" because that would give me the place of 1/128 which would end up being 9, and i need 9 because that is the highest index which i need to place above the sigma symbol. But i dont have the highest index...so how the heck do i get it?

OpenStudy (anonymous):

To find n: As you said, a_n = a_1 r^(n - 1). Because a_n = 1/128, a_1 = 2 and r = 1/2, 1/128 = 2 * (1 / 2)^(n - 1) 1/256 = (1 / 2)^(n - 1) = 2^(1 - n) Take log_2 on both sides. (If you don't have log_2, take log instead and then divide by log(2).) -8 = 1 - n n = 9

OpenStudy (anonymous):

The key there is the log function. Whenever you need to solve for something in the exponent, try log.

OpenStudy (anonymous):

is there any othe way?

OpenStudy (anonymous):

other than the using the log to solve the equation?

OpenStudy (anonymous):

There might be, but I am confident that log is the simplest. One option would be to write a small script that does what you'd do manually, without log.

OpenStudy (anonymous):

ok, for one last concern of mine. Is it correct to use the formula in the problem to find the highest index? Because everyone else didnt seem to care for it

OpenStudy (anonymous):

Of course.

OpenStudy (anonymous):

so it is correct to use the formula?

OpenStudy (anonymous):

When you use the formula, you are clear about your reasoning and you get the right answer, so yes, it is correct.

OpenStudy (anonymous):

lol ok thanks, sorry for the extra question, i failed the test and tommorrow is my last chance to redo it so i want to make sure im certain of everything haha

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

WAIT!'

OpenStudy (anonymous):

in your solution you put that 1/128=2*(1/2)... Then after you multipled the 2 you rewrote as 1/256=...etc. How did you get 1/256?